An ac generator supplies an rms voltage of 140 V and an rms current of 2.0 A at an angular frequency of of 500 rad/s. The generator is connected in series with an inductor, a 200-µF capacitor and an 56-Ω resistor. What is the inductance of the inductor?
Answer #1

we know, Irms = Vrms - 2= Vrms = 7022. Irms Also w= 500 22= R² + (x2-x) = 4900 = 3136+ (col-too? al-wc = 42 LE [42 + 500X200